You paid $957,3 for a 5% 5-year bond, which has a face value of $1000 and pays coupons twice each year. What is the yield-to-maturity? Choose the closest answer

A 2%

B 3%

C 4%

D 5%

E 6%

A bond matured in 15 years has a current yield of 8.35%. The face value is $1000 while the selling price is $1197.93. What is its coupon rate if it pays semi-annual coupon payments?

A 4.18%

B 5.00%

C 8.35%

D 10%

E None of the above
Answer #1

1.

From the given information,

Bond price, PV = 957.3

Coupon, PMT = 0.05*1000/2 = 25

Term, N = 5*2 = 10

Face value, FV = 1000

YTM, R = ?
Using excel function, R = Rate(N,PMT,PV,FV)

Rate(10,-25,957.3,-1000) = 3.00%

Annual YTM = 3*2 = 6%

Option E.

RATE(10,-25,957.3,-1000) 3.00%

2.

Current yield = Annual coupon/Price

8.35% = Annual coupon/1197.93

Annual coupon = 8.35%*1197.93 = 100.03

Coupon rate = Annual coupon/Face value = 100.03/1000 = 10.003%

Option D.